Given m/(1+i)=1-ni, m and n are real numbers, what is the value of m+ni?

Answer by ptaylor(2198)

Left side:
m/(1+i); multiply numerator and denominator by (1-i) and we get:
m(1-i)/(1+1)=m(1-i)/2
Now we have:
m(1-i)/2=1-ni multiply each side by 2
m-mi=2-2ni add mi and also ni to each side
m-mi+mi+ni=2-2ni+ni+mi collect like terms
m+ni=2-ni+mi=2-i(n-m)
Hope this helps---ptaylor